Landon can climb a certain hill at a rate this is 0.5 mph slower than his rate coming down the hill. If it takes him 4 hours to climb the hill and 216 minutes to come down the hill, what is his rate coming down?
Answer by richwmiller(17219)

4 hours=240 minutes climb
3.6= 216 minutes descend
x*3.6=4*(x-.5)
x = 5. mph downhill
x-.5=4.5 mph uphill= 18 miles
5*3.6=18 miles